how much water should be added to 1 gallon of pure antifreeze to obtain a solution that is 15% antifreeze?
Answer by richwmiller(17219)

from view of antifreeze
1*1=0.15*(1+x)
1=0.15+0.15*x
1-0.15=0.15*x
0.85=0.15*x
x=5.66666667
check
1*1=0.15*(1+x)
1*1=0.15*(6.66666667)
1=1.0
ok
or from view of water
1*y+0*1=0.85(1+y)
1*y+0=0.85+0.85y
1*y-0.85y=0.85-0
0.15y=0.85
y=0.85/0.15
y=5.66666667
acid view x 5.66666667= water view y 5.66666667
match 5.66666667=5.66666667